OpenStudy (anonymous):

Please Simplify the square root of 49x (cubed) y (squared)

OpenStudy (anonymous):

49x (cubed) = 7*7 * x*x*x y (squared) = y * y so it should be 7 * 7 * x * x * x * y * y :-)